# Decentralized real-time iterations for distributed NMPC

### v2

@@ V-A Implementation Aspects
With respect to communication, the decentralized RTI scheme requires each subsystem to exchange messages with all neighbors in order to execute Step 4 of Algorithm 1. All other steps in ADMM and dSQP can be carried out without communication. As mentioned in Remark 2, this decentralized ADMM implementation is due to the fact that Step 4 is equivalent to an averaging procedure [Boyd2011, Ch. 7]. Essentially, the step requires each subsystem to exchange predicted state trajectories with neighbors, compute an averaged state trajectory, and then to exchange the averaged trajectories with neighbors. Details about the implementation of the averaging procedure in a DMPC context are provided in [Bestler2019, Stomberg2023]. In this context, we comment on the relation between the dual variables $\lambda$ and $\gamma$. Observe that $\lambda$ is implicitly updated by ADMM in (8b), but only appears in the initialization Step 1 of Algorithm 2. Apart from this initialization, $\lambda$ is only needed in the theoretical convergence analysis and not in the implementation of dSQP. This facilitates the averaging step for computing $z^{l+1}$ as outlined in Appendix B, where the explicit computation of $\lambda^{l+1}$ is omitted. ADMM ensures $\gamma_{i}^{l+1}=E_{i}^{\top}\lambda^{l+1}$ for all $i\in\mathcal{S}$ and for all $k,l\in\mathbb{N}_{0}$. This can be used in closed-loop control to warm start $\gamma_{i}^{0}$ in Step 1 of dSQP with $\gamma_{i}^{k_{\mathrm{max}}}$ from the previous NMPC step. Thus, an initialization $\lambda^{0}$ is only required in the first control step. Possible initializations for the first control step that often work well are $\lambda^{0}=0$ or $\lambda^{0}=\lambda^{\star}$, if available.
With respect to the Hessian, the stability analysis holds for the exact Hessian $H_{i}=\nabla_{z_{i}z_{i}}^{2}L_{i}$, because this guarantees local q-linear convergence of dSQP such that the controller meets Assumption 3. However, the exact Hessian may not always be a favorable choice, because $\nabla_{z_{i}z_{i}}^{2}L_{i}$ can be indefinite outside the dSQP convergence region $\mathcal{B}(p^{\star},\varepsilon)$ and because $\nabla_{z_{i}z_{i}}^{2}L_{i}$ must be evaluated in each dSQP iteration. Instead, the Constrained Gauss-Newton (CGN) method, first introduced as the generalized Gauss-Newton method [Bock1983], provides an alternative which is often effective in NMPC [Messerer2021]. Consider a specialized version of NLP (5) with least squares objectives
@@
where the matrices $M_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ are positive definite and the vectors $m_{i}\in\mathbb{R}^{n}_{i}$ are constants for all $i\in\mathcal{S}$. Such objectives commonly occur in NMPC for setpoint stabilization, see Section VI. The CGN method builds QP (6) with the Gauss-Newton (GN) Hessian approximation $H_{i}=B_{i}\doteq M_{i}^{\top}M_{i}$ for all $i\in\mathcal{S}$. Crucially, the matrices $B_{i}$ are constant, positive definite, and can be evaluated offline.
@@
However, stability guarantees for our proposed RTI scheme when using the GN Hessian are yet unavailable. While the CGN method is locally guaranteed to converge q-linearly in some norm [Messerer2021], the convergence is not necessarily q-linear in the Euclidean norm $\|p^{k}-p^{\star}\|$. This impedes a straight-forward stability proof via Lemma 1 as Assumption 3 may not hold.
VI Numerical Results
@@
We consider the setpoint stabilization of coupled inverted pendulums. Each pendulum is attached to a cart and the carts are coupled via springs as shown in Figure 1. Let $q_{i}$ be the cart position and $\varphi_{i}$ be the angular deviation from the upright position for pendulum $i$. The state of pendulum $i$ is $x_{i}=(q_{i},\dot{q}_{i},\varphi_{i},\dot{\varphi}_{i})\in\mathbb{R}^{4}$ and the input is the force $u_{i}=F_{i}\in[-100\,\text{N},100\,\text{N}]$ applied to the cart. The carts are connected in a chain where each cart is coupled to its neighbors via a spring with stiffness $k=0.1\,$N/m. Let $M_{\mathrm{c}}=2\,$kg and $m=0.25\,$kg denote the masses of each cart and pendulum, respectively, let $l=0.2\,$m denote the length of each pendulum, and denote the gravity of earth by $g=9.81\,\text{m}/\text{s}^{2}$. The continuous-time equations of motion of pendulum $i$ read $\ddot{q}_{i}=(u_{i}+\frac{3}{4}mg\sin(\varphi_{i})\cos(\varphi_{i})-\frac{ml}{2}\dot{\varphi}_{i}^{2}\sin(\varphi_{i})+F_{i}^{\mathrm{left}}+F_{i}^{\mathrm{right}})/(M_{\mathrm{c}}+m-\frac{3}{4}m(\cos(\varphi_{i}))^{2})$ and ${\ddot{\phi}_{i}=\frac{3g}{2l}\sin(\varphi_{i})+\frac{3}{2l}\cos(\varphi_{i})\ddot{q}_{i}}$. If applicable, the coupling forces are $F_{i}^{\mathrm{left}}\doteq k(q_{i-1}-q_{i})$ and $F_{i}^{\mathrm{right}}\doteq k(q_{i+1}-q_{i})$. We discretize the continuous-time dynamics using the explicit Runge-Kutta fourth-order method (RK4) with a shooting interval ${h=40}\,$ms and obtain the discrete-time model $f_{i}^{h}:\mathbb{R}^{4}\times\mathbb{R}\times\mathbb{R}^{{n_{i}^{\mathrm{in}}}}\rightarrow\mathbb{R}^{4}$. For this discretization, we not only keep the control $u_{i}$ constant over the integration step, but also the neighboring positions $q_{i-1}$ and $q_{i+1}$ when evaluating the system dynamics. This simplification preserves the coupling structure, i.e., the discrete-time dynamics of each subsystem only depend on the left and right neighbors. However, the trade-off associated with this simplification is that the integration order with respect to the neighboring states reduces to one. A more accurate alternative is the distributed multiple shooting scheme [Savorgnan2011] which expresses the state trajectories as linear combinations of basis functions, e.g., Legendre polynomials, and matches the basis function coefficients between neighboring subsystems.
@@
VI-A Optimal Control Problem Design
We design OCP (1) such that the value function is a Lyapunov function which satisfies inequalities (3). For the sake of reducing the computational burden, our design does not enforce a terminal constraint in the OCP [Limon2006]. Thus, for the pendulum example, all inequality constraints are input constraints. However, if desired, terminal constraints could also be considered. We choose separable quadratic stage costs $\ell_{i}(x_{i},u_{i})=x_{i}^{\top}Q_{i}x_{i}/2+u_{i}^{\top}R_{i}u_{i}/2,$ where we set $R_{i}=0.001$ and take $Q_{i}=\text{diag}(1,10^{-4},10,10^{-4})$ from [Mills2009]. We further design $V_{\text{f},i}(x_{i})={\beta_{2}}x_{i}^{\top}P_{i}x_{i}/2$, where we find $\beta_{2}\geq 1$ and $P_{i}\in\mathbb{R}^{4\times 4}$ as follows. First, we neglect coupling between neighbors, discretize the dynamics with RK4 at sampling interval $\delta=40\,$ms, and linearize at the origin to obtain a model $(A_{i},B_{i})$ with $A_{i}\in\mathbb{R}^{4\times 4}$ and $B_{i}\in\mathbb{R}^{4\times 1}$ for all $i\in\mathcal{S}$. This allows to solve the algebraic Riccati equation
@@ VI-A Optimal Control Problem Design
we increase $\beta_{2}$ until the matrix ${\Delta Q}\doteq\tilde{Q}-Q_{K}$ is positive definite, where $\tilde{Q}\doteq\beta_{2}(P-A_{K}^{\top}PA_{k}){/\mu}$ and where the paramter $\mu$ is in the notation of [Rawlings2019, Sec. 2.5.5]. We set $\mu=1.01$ and obtain $\beta_{2}=1.1$.
Remark 10 (Decentralized $V_{\mathrm{f}}$ design)
@@
The matrices $P_{i}$ and terminal controllers $K_{i}$ are computed by neglecting the coupling of the dynamics and by solving the Riccati equation for each subsystem individually. This simplification is not guaranteed to stabilize coupled linearized systems in general, but the design stabilizes the chain of pendulums considered here. Hence, the matrix ${\Delta Q}$ is positive definite for sufficiently large $\beta_{2}$. A similar design approach is suggested in [Stewart2011, Remark 4]. More elaborate decentralized linear-quadratic control designs can be found in the classic textbook [Siljak1991]. $\square$
VI-B Swing-up Simulation
@@
We consider the swing up of $S=20$ pendulums to analyze the efficacy of the proposed RTI scheme. Our Matlab implementation uses CasADi to evaulate the derivatives in Step 3 of dSQP [Andersson2019]. The quadratic subproblems in Step 3 of ADMM are solved using OSQP with tolerances $\epsilon_{\text{abs}}=\epsilon_{\text{rel}}=\epsilon_{\text{prim}}=\epsilon_{\text{dual}}=10^{-8}$, where the tolerances are in the notation of [Stellato2020]. After the control input is computed in each NMPC step, the centralized system is simulated using RK4 with integration step size equal to the control sampling interval $\delta$ to obtain the system state at the next sampling instant. We select the penalty parameter $\rho$ of ADMM from the set $\{0.1,1,10,100\}$ and choose $\rho=1$.
@@
We compare three test cases with varying initial conditions, OCP designs, and solver settings as summarized in Table I. The pendulums initially rest in the lower equilibrium position $x_{i}(0)=(q_{i}(0),0,\pi,0)$, where the initial cart displacement is given in Table I. The goal is to steer all pendulums to the upright equilibrium position at $x_{i}=0$. The OCP is designed with quadratic weights as described in Subsection VI-A. The time horizon is chosen as $T=0.4\,$s and the scaling factor in the objective is set to $\beta=1$. The further parameters in Table I are the shooting interval $h$ in the OCP, the discrete-time horizon $N=T/h$, the chosen Hessian for QP (6), the maximum number of SQP iterations per NMPC step $k_{\text{max}}$, the maximum number of ADMM iterations per SQP iteration $l_{\text{max}}$, the total number of decision variables in the centralized OCP $n$, the total number of subsystem equality constraints $n_{g}$, the total number of subsystem inequality constraints $n_{h}$, and the number of consensus constraints $n_{c}$. For all cases, the control sampling interval is set to $\delta=40\,$ms. For cases one and two, we choose the exact Hessian if $\nabla_{z_{i}z_{i}}^{2}L_{i}$ is positive definite and otherwise we select the GN Hessian. For case three, we always choose the GN Hessian approximation. A quadratic penalty term with weight $10^{-5}$ is added to the objective (1a) for all state copies to meet Assumption 6 (ii). We initialize dSQP in the first NMPC step at a solution found by ipopt [Wachter2006]. In all subsequent NMPC steps, we initialize dSQP with the solution produced in the previous control step.
@@
Table II summarizes the simulation results. We analyze the averaged closed-loop control performance
$J_{\text{cl}}\doteq{\frac{1}{T_{f}}}\sum_{t=0}^{t_{n}}\sum_{i\in\mathcal{S}}{\delta}\cdot\ell_{i}(x_{i}(t),u_{i}(t)),$
@@
where $T_{f}$ is the duration of each simulation and $t_{n}$ is the number of NMPC steps per simulation. Furthermore, we analyze the dSQP execution time per NMPC step on a desktop computer. We run each test case ten times to account for varying execution times in between runs and we take two different types of time measurements in each NMPC step. In the first type, we measure the total execution time of one call to dSQP, which is summarized in the second and third columns from the left in Table II. This includes running the specified number of SQP and ADMM iterations for all pendulums in series as well as the costly creation and destruction of intermediate data structures. This is due to the prototypical nature of our Matlab implementation and would be avoided in embedded applications. In the second type of time measurement, summarized in columns four to six of Table II, we measure only imperative code blocks that cannot be avoided in an efficient implementation, and we take measurements for each subsystem individually. This includes calls to CasADi for evaluating derivatives, calls to OSQP for updating and solving QPs in ADMM, and computing Steps 45 of ADMM. Column six summarizes the percentage of per-subsystem solve times that were below the sampling interval.
@@
In all test cases, the pendulums reach the upright equilibrium position while satisfying the input constraints. The initial cart displacements in the first test case are less challenging than in the second case. In fact, the increase in $k_{\text{max}}$ from the first to the second test case is necessary for a successful swing up due to the different initial conditions. That is, the optimizer settings of case one do not provide stability in simulation for case two. While the per-subsystem computation time in the first test case is below the sampling interval $\delta$, the computation time of the second test case would not be real-time feasible. [^1] Non-negligible computation times below the sampling interval can be compensated by solving the OCP for the subsequent control input [Findeisen2006]. Therefore, we increase the shooting interval in the third test case to reduce the OCP size and we select the GN Hessian to reduce the time spent evaluating derivatives such that case three requires less iterations per NMPC step for a successful swing up. On the other hand, this reduces performance as can be seen from the prolonged settling time and the increase in $J_{\text{cl}}$. Choosing a shooting interval $h>\delta$ is a well-known technique for reducing RTI computation times [Diehl2002]. However, the stability guarantees provided by Theorem 2 only hold if $h=\delta$ and if the exact Hessian is chosen.
The time measurements, obtained for a prototypical Matlab implementation, show that the per-subsystem computation times are mostly below the control sampling interval. This does not imply that an efficient decentralized implementation would necessarily be real-time feasible, in particular if communication latencies add to the computation time. [^2] Latency measurements for a decentralized dSQP implementation indicate that a worst-case latency of approximately 101010\,10ms can be expected for running $k_{\text{max}}=1$ SQP and $l_{\text{max}}=6$ ADMM iterations in a laboratory setting [Stomberg2023]. However, the obtained computation times indicate that control sampling intervals in the millisecond range are conceivable via the decentralized RTI scheme.
TABLE I: Test case specifications. All cases consider a sampling interval $\delta=40\,$ms.
Case $q_{i}(0)$ $h\,$[ms] $N$ $H$ $k_{\text{max}}$ $l_{\text{max}}$ $n$ $n_{g}$ $n_{h}$ $n_{c}$
1 $-1^{i}$ 40 10 exact 1 6 1518 880 440 418

### v3

@@ 5.1 Implementation Aspects
With respect to communication, the decentralized RTI scheme requires each subsystem to exchange messages with all neighbors in order to execute Step 4 of Algorithm 1. All other steps in ADMM and dSQP can be carried out without communication. As mentioned in Remark 2, this decentralized ADMM implementation is due to the fact that Step 4 is equivalent to an averaging procedure [34, Ch. 7]. Essentially, the step requires each subsystem to exchange predicted state trajectories with neighbors, compute an averaged state trajectory, and then to exchange the averaged trajectories with neighbors. Details about the implementation of the averaging procedure in a DMPC context are provided in [16, 29]. In this context, we comment on the relation between the dual variables $\lambda$ and $\gamma$. Observe that $\lambda$ is implicitly updated by ADMM in (8b), but only appears in the initialization Step 1 of Algorithm 2. Apart from this initialization, $\lambda$ is only needed in the theoretical convergence analysis and not in the implementation of dSQP. This facilitates the averaging step for computing $z^{l+1}$ as outlined in Appendix 9, where the explicit computation of $\lambda^{l+1}$ is omitted. ADMM ensures $\gamma_{i}^{l+1}=E_{i}^{\top}\lambda^{l+1}$ for all $i\in\mathcal{S}$ and for all $k,l\in\mathbb{N}_{0}$. This can be used in closed-loop control to warm start $\gamma_{i}^{0}$ in Step 1 of dSQP with $\gamma_{i}^{k_{\mathrm{max}}}$ from the previous NMPC step. Thus, an initialization $\lambda^{0}$ is only required in the first control step. Possible initializations for the first control step that often work well are $\lambda^{0}=0$ or $\lambda^{0}=\lambda^{\star}$, if available.
With respect to the Hessian, the stability analysis holds for the exact Hessian $H_{i}=\nabla_{z_{i}z_{i}}^{2}L_{i}$, because this guarantees local q-linear convergence of dSQP such that the controller meets Assumption 3. However, the exact Hessian may not always be a favorable choice, because $\nabla_{z_{i}z_{i}}^{2}L_{i}$ can be indefinite outside the dSQP convergence region $\mathcal{B}(p^{\star},\varepsilon)$ and because $\nabla_{z_{i}z_{i}}^{2}L_{i}$ must be evaluated in each dSQP iteration. Instead, the Constrained Gauss-Newton (CGN) method, first introduced as the generalized Gauss-Newton method [43], provides an alternative which is often effective in NMPC [44]. Consider a specialized version of NLP (5) with least squares objectives
@@
where the matrices $M_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ are positive definite and the vectors $m_{i}\in\mathbb{R}^{n_{i}}$ are constants for all $i\in\mathcal{S}$. Such objectives commonly occur in NMPC for setpoint stabilization, see Section 6. The CGN method builds QP (6) with the Gauss-Newton (GN) Hessian approximation $H_{i}=B_{i}\doteq M_{i}^{\top}M_{i}$ for all $i\in\mathcal{S}$. Crucially, the matrices $B_{i}$ are constant, positive definite, and can be evaluated offline.
@@
However, stability guarantees for our proposed RTI scheme when using the GN Hessian are yet unavailable. While the CGN method is locally guaranteed to converge q-linearly in some norm [44], the convergence is not necessarily q-linear in the Euclidean norm $\|p^{k}-p^{\star}\|$. This impedes a straight-forward stability proof via Lemma 1 as Assumption 3 may not hold.
6 Numerical Results
@@
We consider the setpoint stabilization of coupled inverted pendulums. Each pendulum is attached to a cart and the carts are coupled via springs as shown in Figure 1. Let $q_{i}$ be the cart position and $\varphi_{i}$ be the angular deviation from the upright position for pendulum $i$. The state of pendulum $i$ is $x_{i}=(q_{i},\dot{q}_{i},\varphi_{i},\dot{\varphi}_{i})\in\mathbb{R}^{4}$ and the input is the force $u_{i}=F_{i}\in[-100\,\text{N},100\,\text{N}]$ applied to the cart. The carts are connected in a chain where each cart is coupled to its neighbors via a spring with stiffness k=0.1k=0.1\,N/m. Let $M_{\mathrm{c}}=2\,$kg and m=0.25m=0.25\,kg denote the masses of each cart and pendulum, respectively, let l=0.2l=0.2\,m denote the length of each pendulum, and denote the gravity of earth by $g=9.81\,\text{m}/\text{s}^{2}$. The continuous-time equations of motion of pendulum $i$ read $\ddot{q}_{i}=(u_{i}+\frac{3}{4}mg\sin(\varphi_{i})\cos(\varphi_{i})-\frac{ml}{2}\dot{\varphi}_{i}^{2}\sin(\varphi_{i})+F_{i}^{\mathrm{left}}+F_{i}^{\mathrm{right}})/(M_{\mathrm{c}}+m-\frac{3}{4}m(\cos(\varphi_{i}))^{2})$ and $\ddot{\phi}_{i}=\frac{3g}{2l}\sin(\varphi_{i})+\frac{3}{2l}\cos(\varphi_{i})\ddot{q}_{i}$}. If applicable, the coupling forces are $F_{i}^{\mathrm{left}}\doteq k(q_{i-1}-q_{i})$ and $F_{i}^{\mathrm{right}}\doteq k(q_{i+1}-q_{i})$. We discretize the continuous-time dynamics using the explicit Runge-Kutta fourth-order method (RK4) with a shooting interval h=40{h=40}\,ms and obtain the discrete-time model $f_{i}^{h}:\mathbb{R}^{4}\times\mathbb{R}\times\mathbb{R}^{n_{i}^{\mathrm{in}}}\rightarrow\mathbb{R}^{4}$. For this discretization, we not only keep the control $u_{i}$ constant over the integration step, but also the neighboring positions $q_{i-1}$ and $q_{i+1}$ when evaluating the system dynamics. This simplification preserves the coupling structure, i.e., the discrete-time dynamics of each subsystem only depend on the left and right neighbors. However, the trade-off associated with this simplification is that the integration order with respect to the neighboring states reduces to one. A more accurate alternative is the distributed multiple shooting scheme [45] which expresses the state trajectories as linear combinations of basis functions, e.g., Legendre polynomials, and matches the basis function coefficients between neighboring subsystems.
@@
6.1 Optimal Control Problem Design
We design OCP (1) such that the value function is a Lyapunov function which satisfies inequalities (3). For the sake of reducing the computational burden, our design does not enforce a terminal constraint in the OCP [32]. Thus, for the pendulum example, all inequality constraints are input constraints. However, if desired, terminal constraints could also be considered. We choose separable quadratic stage costs $\ell_{i}(x_{i},u_{i})=x_{i}^{\top}Q_{i}x_{i}/2+u_{i}^{\top}R_{i}u_{i}/2,$ where we set $R_{i}=0.001$ and take $Q_{i}=\text{diag}(1,10^{-4},10,10^{-4})$ from [46]. We further design $V_{\text{f},i}(x_{i})={\beta_{2}}x_{i}^{\top}P_{i}x_{i}/2$, where we find $\beta_{2}\geq 1$ and $P_{i}\in\mathbb{R}^{4\times 4}$ as follows. First, we neglect coupling between neighbors, discretize the dynamics with RK4 at sampling interval $\delta=40\,$ms, and linearize at the origin to obtain a model $(A_{i},B_{i})$ with $A_{i}\in\mathbb{R}^{4\times 4}$ and $B_{i}\in\mathbb{R}^{4\times 1}$ for all $i\in\mathcal{S}$. This allows to solve the algebraic Riccati equation
@@ 6.1 Optimal Control Problem Design
we increase $\beta_{2}$ until the matrix $\Delta Q\doteq\tilde{Q}-Q_{K}$ is positive definite, where $\tilde{Q}\doteq\beta_{2}(P-A_{K}^{\top}PA_{k})/\mu$ and where the paramter $\mu$ is in the notation of [42, Sec. 2.5.5]. We set $\mu=1.01$ and obtain $\beta_{2}=1.1$.
Remark 18 (Decentralized $V_{\mathrm{f}}$ design).
@@
The matrices $P_{i}$ and terminal controllers $K_{i}$ are computed by neglecting the coupling of the dynamics and by solving the Riccati equation for each subsystem individually. This simplification is not guaranteed to stabilize coupled linearized systems in general, but the design stabilizes the chain of pendulums considered here. Hence, the matrix $\Delta Q$ is positive definite for sufficiently large $\beta_{2}$. A similar design approach is suggested in [17, Remark 4]. More elaborate decentralized linear-quadratic control designs can be found in the classic textbook [3]. $\square$
6.2 Swing-up Simulation
@@
We consider the swing up of $S=20$ pendulums to analyze the efficacy of the proposed RTI scheme. Our Matlab implementation is available online and uses CasADi to evaulate the derivatives in Step 3 of dSQP [47]. [^1] https://github.com/OptCon/real-time-dmpc The quadratic subproblems in Step 3 of ADMM are solved using OSQP with tolerances $\epsilon_{\text{abs}}=\epsilon_{\text{rel}}=\epsilon_{\text{prim}}=\epsilon_{\text{dual}}=10^{-8}$, where the tolerances are in the notation of [48]. After the control input is computed in each NMPC step, the centralized system is simulated using RK4 with integration step size equal to the control sampling interval $\delta$ to obtain the system state at the next sampling instant. We select the penalty parameter $\rho$ of ADMM from the set $\{0.1,1,10,100\}$ and choose $\rho=1$.
@@
We compare three test cases with varying initial conditions, OCP designs, and solver settings as summarized in Table 1. The pendulums initially rest in the lower equilibrium position $x_{i}(0)=(q_{i}(0),0,\pi,0)$, where the initial cart displacement is given in Table 1. The goal is to steer all pendulums to the upright equilibrium position at $x_{i}=0$. The OCP is designed with quadratic weights as described in Subsection 6.1. The time horizon is chosen as T=0.4T=0.4\,s and the scaling factor in the objective is set to $\beta=1$. The further parameters in Table 1 are the shooting interval $h$ in the OCP, the discrete-time horizon $N=T/h$, the chosen Hessian for QP (6), the maximum number of SQP iterations per NMPC step $k_{\text{max}}$, the maximum number of ADMM iterations per SQP iteration $l_{\text{max}}$, the total number of decision variables in the centralized OCP $n$, the total number of subsystem equality constraints $n_{g}$, the total number of subsystem inequality constraints $n_{h}$, and the number of consensus constraints $n_{c}$. For all cases, the control sampling interval is set to $\delta=40\,$ms. For cases one and two, we choose the exact Hessian if $\nabla_{z_{i}z_{i}}^{2}L_{i}$ is positive definite and otherwise we select the GN Hessian. For case three, we always choose the GN Hessian approximation. A quadratic penalty term with weight $10^{-5}$ is added to the objective (1a) for all state copies to meet Assumption 6 (ii). We initialize dSQP in the first NMPC step at a solution found by IPOPT [49]. In all subsequent NMPC steps, we initialize dSQP with the solution produced in the previous control step.
@@
Table 2 summarizes the simulation results. We analyze the averaged closed-loop control performance
$J_{\text{cl}}\doteq\frac{1}{t_{n}+1}\sum_{t=0}^{t_{n}}\sum_{i\in\mathcal{S}}\ell_{i}(x_{i}(t),u_{i}(t)),$
@@
where $t_{n}\doteq T_{f}/\delta$ and where $T_{f}=10\,$s is the simulated time span. Furthermore, we analyze the dSQP execution time per NMPC step on a desktop computer. We run each test case ten times to account for varying execution times in between runs and we take two different types of time measurements in each NMPC step. In the first type, we measure the total execution time of one call to dSQP, which is summarized in the second and third columns from the left in Table 2. This includes running the specified number of SQP and ADMM iterations for all pendulums in series as well as the costly creation and destruction of intermediate data structures. This is due to the prototypical nature of our Matlab implementation and would be avoided in embedded applications. In the second type of time measurement, summarized in columns four to six of Table 2, we measure only imperative code blocks that cannot be avoided in an efficient implementation, and we take measurements for each subsystem individually. This includes calls to CasADi for evaluating derivatives, calls to OSQP for updating and solving QPs in ADMM, and computing Steps 45 of ADMM. Column six summarizes the percentage of per-subsystem solve times that were below the sampling interval.
@@
In all test cases, the pendulums reach the upright equilibrium position while satisfying the input constraints. The initial cart displacements in the first test case are less challenging than in the second case. In fact, the increase in $k_{\text{max}}$ from the first to the second test case is necessary for a successful swing up due to the different initial conditions. That is, the optimizer settings of case one do not provide stability in simulation for case two. While the per-subsystem computation time in the first test case is below the sampling interval $\delta$, the computation time of the second test case would not be real-time feasible. [^2] Non-negligible computation times below the sampling interval can be compensated by solving the OCP for the subsequent control input [50]. Therefore, we increase the shooting interval in the third test case to reduce the OCP size and we select the GN Hessian to reduce the time spent evaluating derivatives such that case three requires less iterations per NMPC step for a successful swing up. On the other hand, this reduces performance as can be seen from the prolonged settling time and the increase in $J_{\text{cl}}$.
Remark 19 (Stability, shooting interval, and GN Hessian).
Test case three selects a larger shooting interval and the GN Hessian to accelerate the computations. In general, choosing a shooting interval $h>\delta$ is a well-known technique for reducing RTI computation times [4] and we found the GN Hessian to work well in hardware experiments [29, 30]. However, while we observe stability for the pendulum simulation, the guarantees provided by Theorem 12 only hold if $h=\delta$ and if the exact Hessian $H\doteq\nabla_{zz}L$ is chosen. $\square$
The time measurements, obtained for a prototypical Matlab implementation, show that the per-subsystem computation times are mostly below the control sampling interval. This does not imply that an efficient decentralized implementation would necessarily be real-time feasible, in particular if communication latencies add to the computation time. [^3] Latency measurements for a decentralized dSQP implementation indicate that a worst-case latency of approximately 1010\,ms can be expected for running $k_{\text{max}}=1$ SQP and $l_{\text{max}}=6$ ADMM iterations in a laboratory setting [29]. However, the obtained computation times indicate that control sampling intervals in the millisecond range are conceivable via the decentralized RTI scheme.
Table 1: Test case specifications. All cases consider a sampling interval $\delta=40\,$ms.
Case $q_{i}(0)$ hh\,[ms] NN $H$ $k_{\text{max}}$ $l_{\text{max}}$ $n$ $n_{g}$ $n_{h}$ $n_{c}$
1 $1^{i}$ 40 10 exact 1 6 1518 880 440 418